### v1

# S-EQA: Tackling Situational Queries in Embodied Question Answering

Abstract
We present and tackle the problem of Embodied Question Answering (EQA) with Situational Queries (S-EQA) in a household environment. Unlike prior EQA work tackling simple queries that directly reference target objects and quantifiable properties pertaining them, EQA with situational queries (such as “Is the bathroom clean and dry?”) is more challenging, as the agent needs to figure out not just what the target objects pertaining to the query are, but also requires a consensus on their states to be answerable. Towards this objective, we first introduce a novel Prompt-Generate-Evaluate (PGE) scheme that wraps around an LLM’s output to create a dataset of unique situational queries, corresponding consensus object information, and predicted answers. PGE maintains uniqueness among the generated queries, using multiple forms of semantic similarity. We validate the generated dataset via a large scale user-study conducted on M-Turk, and introduce it as S-EQA, the first dataset tackling EQA with situational queries. Our user study establishes the authenticity of S-EQA with a high of the generated queries being deemed answerable, given the consensus object data. Conversely, we observe a low correlation of on the LLM-predicted answers to human-evaluated ones; indicating the LLM’s poor capability in directly answering situational queries, while establishing S-EQA’s usability in providing a human-validated consensus for an indirect solution. We evaluate S-EQA via Visual Question Answering (VQA) on VirtualHome, which unlike other simulators, contains several objects with modifiable states that also visually appear different upon modification — enabling us to set a quantitative benchmark for S-EQA. To the best of our knowledge, this is the first work to introduce EQA with situational queries, and also the first to use a generative approach for query creation. Our work introduces a new task and approach that aims to improve the usability of embodied agents in real-world household environments.
I Introduction
Embodied Question Answering (EQA) introduced by Das. et. al., [1] involves an agent exploring its surroundings to find an answer to a query. Prior work on this task [2, 3, 4, 5] has primarily dealt with what we term “simple queries”, that are directly answerable from states or properties of objects from the simulator. For instance, questions like “What is the color of the sofa?” or “Is there an object nearby to help me cut vegetables?” are directly related to an object present in the environment, requiring a mix of vision-language grounding with elementary commonsense to solve the task. While important, this type of question-answering capability does not add much utility to a real-world household robot, as the user asking questions is likely to already know answers to simple questions about their surroundings.
| Dataset | Navigation + QA | Multi-Target | Interaction | Abstract Queries | Object State Consensus | Query Creation |
| EQA [1] | ✓ | - | - | - | - | Rule-Based |
| MT-EQA [4] | ✓ | ✓ | - | - | - | Rule-Based |
| IQA [5] | ✓ | ✓ | ✓ | - | - | Rule-Based |
| K-EQA [6] | ✓ | ✓ | - | ✓ | - | Rule-Based |
| S-EQA (Ours) | ✓ | ✓ | - | ✓ | ✓ | Generative |
In contrast, answering queries such as “Is the kitchen ready for meal preparation?” requires a sense of situational awareness in gauging the status of multiple objects in the environment. While not uniformly defined, a consensus does exist on standard conditions that can be used to answer such questions. This refers to a set of generally accepted object states and conditions (such as a Knife being Present) for successfully evaluating the query. A household robot deployed in the real world is likely to encounter such queries that require it to utilize consensus knowledge pertaining which objects and associated states are commonly associated with the situation. We term these as situational queries, and in this work, tackle the twofold challenge of 1) conjuring authentic situational queries and corresponding consensus object data and 2) highlighting the role of the consensus in simplifying the process of responding to situational queries.
A query such as “Is the house ready for sleeptime?” is a situational query, since there exist generally accepted object states, such as dimmed lights, closed windows, locked doors, etc., which most people associate with a house being ready for nighttime. However, a query like “Can I bring guests over?”, while situational, is too ambiguous, as the answer relies on various unstated factors and user preferences on household norms, and event-specific constraints. For situational queries to be validatable, it is important for them to not be too ambiguous so as to allow for a consensus to exist. Our work places emphasis on this, and follows an in-context learning scheme on an LLM to avoid overly simplistic nor excessively ambiguous queries.
Developing a dataset for EQA in the past has been limited to constructing simple queries using entities from a simulator [4, 5], with some work extending it to abstract queries built on scene-graphs [6]. Posing situational queries using these existing techniques is however nontrivial, as they stem from human intuition and contextual awareness of the environment. Setting up real-world experiments to collect this data is also cumbersome; if not having an intrusive agent capturing conversations at home, an offline approach would be cognitively heavy, requiring humans to come up with queries for imaginary situations. As such, in our work, we use an LLM to generate and study such queries, while ensuring they cover unique concepts stemming from various situations. Table I compares our dataset with existing ones.
Main Results: To tackle these issues, we leverage the commonsense capabilities of Large Language Models (LLMs) and present an approach to generate situational queries and associated consensus object states and relationships. LLMs allow us to generate situational query data for a variety of scenarios which are nontrivial to obtain via real-world data collection. The main contributions of our work are as follows:-
We propose a novel Large Language Model (LLM)-based method for generating Situational Queries for Embodied Question Answering (EQA). Unlike simple queries that directly reference particular objects, situational queries necessitate an understanding of the consensus objects and states pertaining a situation.
Our proposed Prompt-Generate-Evaluate (PGE) scheme wraps around an LLM to generate a wide array of unique and challenging queries covering various household scenarios. Along with the queries, we also generate corresponding consensus object information, pointing to object states and relationships relevant to the queries.
The generated queries are validated via a large scale MTurk [7] user study to produce S-EQA, a LLM-generated dataset containing situational queries, associated consensus object data, and predicted answers. While the number of queries is small due to their highly abstract nature, PGE ensures diversity in the queries across rooms.
We quantitatively evaluate S-EQA by performing Visual Question Answering (VQA) on VirtualHome images with both situational queries (Room VQA) and consensus data (Object VQA) to establish a benchmark.
We observe that 97.26% of users find the generated queries to be answerable provided the consensus data, while only 46.2% of user answers align with the LLM predicted ones. From the former, we can infer the authenticity of LLM generated data in producing human-like consensus states, while the latter indicates the LLM cannot be used to directly answer situational queries; establishing the need and use-case for S-EQA, a human-validated dataset that provides situational queries along with consensus objects and states.
II Related Work
Embodied Question Answering (EQA): EQA is a popular task [8, 5, 9] that involves a robot attempting to answer a query by exploring and gathering information from its surroundings. It was introduced by Das et. al [1] as a single-target question-answering dataset. Yu et. al. [4] further extended this task to a multi-object setting, MT-EQA, where correspondences between multiple target objects present in the question itself need to be analyzed before arriving at an answer. Both these works use a template-based approach to generating questions, and this was deemed sufficient since the queries are objective in nature mentioning the objects of interest in their query itself. This however is a limitation, since a majority of queries tend to be subjective or situational in nature [10], especially in a household environment.
Sintean et. al in K-EQA [6] attempt to tackle this problem, by using scene-graph data and formal grammar to postulate queries. However, these queries are still limited to simple object grounding within a scene, and do not require any form of situational awareness about the readiness of an environment. For instance, S-EQA considers queries like “Is the living room ready for a movie night?”, and presents consensus multiple object states (TV: On, Popcorn: Cooked, On Table, etc.) to measure this readiness. For the same example, a K-EQA query would have been constructed as “Is there a device for watching movies in the living room?”, which requires only scene knowledge for an answer, but does not require situational knowledge about the consensus states of objects for a movie night.
In our work, we utilize an LLM to generate situational queries and approximate human consensus on object states required to answer the query. Our queries are an abstraction above the scene-knowledge based K-EQA queries, requiring consensus knowledge about multiple object states for validation.
LLMs for Query Generation: A recent trend has seen the emergence of Large Language Models (LLMs) such as GPT-3.5 [11], GPT-4 [12], LLaMA [13], Palm [14] and Palm-e [15]. These models are well known for their commonsense reasoning capabilities [16], especially in the context of question-answering. In embodied AI, these capabilities have recently been leveraged to produce superior planning and decision-making strategies during navigation [17, 18, 19, 20]. In our case, we wish to generate situational queries for embodied question answering, which as defined in Section I require consensus about object states and relationships in the environment. As such, we are motivated to use LLMs for generating these queries and ask if their commonsense reasoning capabilities can sufficiently represent the consensus needed to answer situational queries.
III Situational EQA: Generation, Analysis, and Evaluation
Having motivated the problem of Situational EQA, we now present a detailed description of our approach. We present three phases — generating situational data using LLMs, validating this data via segregation and a user survey, and finally evaluating it on the VirtualHome simulator. We hope the generative approach being used to develop S-EQA helps foster future research in this domain.
III-A Situational Query Definition
Situational queries in our work are defined as questions requiring the assessment of a collection of consensus objects and states to reach an existential ‘Yes/No’ answer. For example, a query such as “Is the livingroom ready for movie night?” has a Yes/No answer based on whether consensus objects like the TV, lights, etc. are in their appropriate state. Additionally, we want to ensure that the queries are related to a situation and are not simple in nature, i.e., the target objects and states are not in the query itself. For example, “Is the sofa blue?” [1] or “How many objects can be used to hold rice?” [6] are simple queries, as they do not deal with a situation, but rather pose a query referencing objects in the environment. In contrast, “Is the dining area set up for dinner?” is situational, as it doesn’t explicitly mention any objects, but requires the identification of relevant objects (eg. table, napkins), and a consensus on what their states should be (open, present) for the answer. Considering the LLM as a function that generates queries , and to be a set of all possible queries generated by , we want situational queries , which satisfies the following conditions -
The query must not directly reference any object, i.e., it must omit words from a set which includes “object”, “table”, “chair”, etc.
The query must be a question.
The answer to the query must be binary i.e., Yes/No, and not subjective in nature.
Formally, we can define as:
This formulation ensures that consists only of queries that are abstract, questioning in form, and expect a binary Yes/No response. We present these conditions as in-context rules to the LLM to curate it’s responses.
III-B Generating Situational Queries: S-EQA
Figure 2 presents an overview of our generation approach. We generate datapoints of the format - [Situational Query, Object-States, Object-Relationships, LLM Answer]. A sample datapoint is as follows -
Query: “Is the bathroom ready for a shower?”
Object-States : [lightswitch: [‘On’], towels: [‘Present’], Soap: [‘Present’] …]
Object-Relationships : [lightswitch inside bathroom, towels inside bathroom, … ]
LLM Answer: ‘Yes’
Here, the LLM points to a set of object states and relationships (such as ‘towels inside bathroom being present’) that it believes need to be checked to determine if the bathroom is ready for a shower. We also predict an answer to the query that is later verified in our user study.
III-B1 VirtualHome Setup
We use the VirtualHome [21] simulator for defining ground truth objects, states and properties in our experiments. It contains 7 different household environments containing over 390 objects. Unlike other embodied simulators, VirtualHome is unique in that it contains several household objects including interactable ones, with easily modifiable locations and states. The objects are also visually different upon modification, and this allows us to perform VQA later on to evaluate our dataset.
All objects in VirtualHome have a set of state properties, which are either physical Open/Closed (eg. for a Door), and/or On/Off (such as lights) or Present/None (such as Apple). These objects also have a relationship with either the room or another object. For example, Apple INSIDE Fridge and Fridge INSIDE livingroom are both valid relationships. The simulator has 4 different rooms into which all objects are placed - [livingroom, bedroom, kitchen, bathroom]. We use these room markings to group and analyze the generated queries later on.
III-B2 Prompt-Generate-Evaluate (PGE) Scheme
Figure 2 presents an overview of our Prompt-Generate-Evaluate scheme for generating data. We perform analytical operations on the LLM’s generated datapoints to ensure they fit into our definition of situational queries, while also maintaining diversity. The System Prompt for our LLM contains a list of all object states and relationships. We also provide additional rules and in-context examples to improve the generation of prompts. For instance, explicitly mentioning “Do not conjure any new object states” ensures that the generated datapoints do not contain states that are not achievable in VirtualHome. Additionally, we also provide a set of “Generated Queries” as feedback to the system prompt. These represent queries that the LLM has already generated, and in the User prompt, encourage the LLM to generate queries different from these. Our strategy for choosing these queries is discussed in the feedback measures presented below.
Our User Prompt asks to generate datapoints in the format specified earlier, and provides a few sample datapoints. We also provide some in-context rules to ensure that the generated queries are not too simplistic, not too ambiguous, and are also unique. For instance, we explicitly state ‘The situational query must not reference any object or even contain the word ‘object’.’ and ‘The query must have a Yes/No answer.’. Additionally, we ask it to generate queries very different from what is already present in the Generated Queries in the System Prompt, to avoid simple sentence rephrasing.
To ensure that our Situational Query Database contains data points that adhere to the specifications presented above, we take the following two feedback measures:
After generating data points, we embed the situational queries using RoBERTa [22] to get sentence embeddings . We perform agglomerative clustering [23] on to accumulate them into groups of queries based on sentence similarity. We then pick the query closest to the centroid of each of these clusters to get queries that best represent the diversity in the generated data. Finally, we update the system prompt with these queries.
Simultaneously, we compute the cosine similarity of the generated query embeddings with the queries already present in the Situational Query Database . If is greater than a particular threshold (), we continue the “conversation” with the LLM, by presenting it with a re-generation prompt. We mention the percentage of similarity, and ask the LLM to re-generate a new batch of datapoints.
These feedback measures ensure contains unique situational queries that are not simple rewordings, while also encouraging diversity in the generated situations.
The total number of datapoints generated is always less than or equal to the product of the looping parameters and , i.e., . While our approach allows for the generation of potentially infinite datapoints, we observe a saturation point where the LLM starts to generate similar-looking queries (“Is the bathroom clean and dry?” vs “Has the bathroom been cleaned today?”). This is indeed a drawback of such a generative approach, and can be attributed to the abstract nature of the queries being generated, which do not pertain a particular object, but related to a household situation. Therefore, to effectively assess the utility of the data, we cut off generation at data points. We validate this data via a large-scale MTurk user study, and present this as S-EQA, a dataset containing diverse and challenging situational queries along with consensus object data to help answer them.
III-C Validating S-EQA: t-SNE Plots and MTurk Survey
Our objective with validating S-EQA is not just to measure the authenticity of the generated data, but also to establish necessity and the real-world usability of our dataset. First, we first visualize the queries in using t-SNE plots to analyze their diversity. We then conduct a Mechanical Turk based user study to validate the generated data.
III-C1 Analyzing S-EQA with t-SNE
Figure 3 illustrates t-SNE plots we derive from . These are as follows -
Room Categorization: Using the room properties obtained from VirtualHome (section III-B1), we group the questions into 6 categories - [kitchen, livingroom, bedroom, bathroom, multi-room, no room] (Ref. Fig. 3(a)). We infer from this plot that most questions being generated belong to one particular room, as opposed to multiple ones. This suggests that the LLM tends to generate situational queries that are usually room-specific, referring to multiple objects in a single area. As such, multi-room scenarios (like “Is the house ready for sleeptime?”) are less common in the generated data.
Situational Categorization: Here, we use the LLM as a classifier to predict if the generated questions are situational according to our definition in III-A. We prompt and present in-context examples accordingly. We notice that of generated questions (Ref. Fig. 3(b)) are classified as situational according to the LLM. As the LLM classifications have several false positives, this result motivates us to use human-validation to authenticate the dataset.
Spatio-Temporal Categorization: Here, we use the LLM to categorize the generated queries as spatial or temporal. Temporal questions rely on time-sensitive knowledge about a dynamic environment (Refer Fig. 3(c)). For instance, a query like “Was someone sleeping in the bedroom today?” is temporal, as it requires knowledge of past environmental states to be answerable. Conversely, a query like “Is the kitchen ready for cooking?” is spatial, as it requires the agent to gauge only current states. We observe that of the generated questions are Spatial ones, showing that PGE prefers modeling spatial relations over temporal ones.
III-C2 Assessing S-EQA’s Utility with MTurk
We conduct a large-scale Mechanical Turk [7] study to gauge S-EQA’s authenticity and usability in answering real-world queries.
In our first experiment, we assess whether the consensus object data being generated adequately facilitates answering situational queries. Workers are shown a query, images of the corresponding room, and images of objects set to consensus states, while the LLM Answer remains hidden to negate bias. Given options of Yes, No, and Cannot Answer, a worker’s capacity to give a definitive response, i.e., choosing Yes or No, serves as an indicator of whether the generated data is in alignment with general user consensus. For instance, the query “Is the house ready for sleeptime?”, if unanimously agreed upon by workers when shown images of dimly-lit rooms, closed doors, and windows, validates the data point.
Subsequently, in a second experiment, workers are presented with an image of a room accompanied by a situational S-EQA query (For eg. Is someone working in the bedroom? with top-views of the bedroom) and, separately, images of consensus objects paired with simpler, objective queries about their state (For eg., Is the Computer On? with an image of a computer being on). For the simple query annotation, we additionally sample ‘Yes/No’ questions the SimVQA [24] dataset to even out the number of situational and simple queries. An improved annotator performance in simple over situational queries would indicate S-EQA’s real-world usability in providing human-authenticated consensus object-states for situational queries which are relatively simpler to answer.
III-D Evaluating S-EQA in VirtualHome
Assuming low-level navigation to be solved, we quantitatively evaluate S-EQA via Visual Question Answering (VQA). VirtualHome provides us with a household environment where object states visually change upon modifying states. We present VQA models with the same setup as described in our first MTurk experiment presented in III-C2, and compare the results. Figure 4 presents an overview.
Room VQA (Situational Query): Here, we present the VQA model with multiple top-view images of a relevant room as input, and ask it to answer the associated situational query with Yes/No/Cannot Answer. Most of our queries are associated with one particular room (see Figure 3(a)), which is often part of the query itself. During our experimentation, we observed a clear bias in model performance with such queries. For instance, if the query contained ‘living room’, and the image was that of the living room, the answer would always be yes, irrespective of the actual context of the query. To mitigate this, we replace any mention of a room in the query with the generic term - “this place”.
Object VQA (Simple Query): Here, we use consensus object states linked with the situational query to formulate simple objective questions for a VQA model. These are of the form -
Is the OBJECT STATE?
We provide the VQA model with questions along with images of the consensus objects in their designated states.
For both Room VQA and Object VQA, we measure the accuracy and F1-scores of prediction across all the images, using the human-verified answers as ground truth. We use a threshold of on the accuracy and F1 scores to determine if the query has successfully been answered. Additionally, we also compute the Joint Accuracy (), and Joint F1 () scores which is a success case when either of the Room or Object VQA are successful. Formally, let be the accuracy or F1 result for Room VQA and be the same result for Object VQA, where . indicates a match with the ground truth for Room VQA, and similarly, indicates a match for Object VQA.
is then defined as:
and give us an upper bound on the performance of VQA models on S-EQA in VirtualHome. For , we set a threshold of above which the score would be deemed satisfactory.
IV Results and Inferences
We use GPT-4 [12] as the LLM of choice to generate situational queries in our experiments. Our ablations with GPT-3.5 showed poor quality in terms of generation, with many outputs not fitting into our specified datapoint format.
IV-A Preparing the Simulator
We first program the object states and relationships from the generated data points into the VirtualHome environment. This also includes filtering out a small subset of bad states which are not feasible. For instance, ‘computer INSIDE fridge’ is not a valid relationship for us to set, and the simulator prevents this.
For VQA and the MTurk user study, we require images of the room and corresponding consensus objects set to their appropriate states. We gather room images by considering the 4 corners of each room, with a camera pointing downwards towards the center. Choosing images in this way allows us to achieve maximum coverage of objects present.
Then for object images, we use the first-person view of a human placed in the environment. The human is instructed to walk to each of the objects present in a S-EQA datapoint, and we grab an image when they are present in front of them. This approach for capturing images proved simpler to implement and is in-line with the visual perspective of an agent grabbing an object.
IV-B Quantitative Results and Benchmark
We use BLIP-2 [25], BLIP [26] and ViLT [27], all finetuned on VQAv2 [28] to evaluate S-EQA, and these results are presented in Table II.
| VQA Model | Accuracy % | F1-Score % | |
| Room | BLIP-2 [25] | 42.31 | 52.51 |
| BLIP [26] | |||
| ViLT [27] | |||
| Object | BLIP-2 [25] | 58.26 | 68.75 |
| BLIP [26] | |||
| ViLT [27] | |||
| Joint | BLIP-2 [25] | 89.6 | |
| BLIP [26] | 83.75 | ||
| ViLT [27] |
Observe the higher Object-VQA accuracies and F1 scores over Room-VQA across each approach. This suggests that answering situational queries by looking at images at a room level is significantly harder than answering simple queries about the states of target objects. Our second MTurk experiment (Refer III-C2) further corroborates this result, suggesting S-EQA’s usability in providing the simpler consensus object data for answering situational queries.
The high values on joint scores indicate an upper bound on performance. While still better than RoomVQA, ObjectVQA models have mediocre performance overall, despite the queries being simple. This can be attributed to inference on synthetic images from VirtualHome, which these models have not been trained on. Our real-world experiments suggest improved performance, confirming this domain gap.
Figure 5 provides a distribution of the query lengths in S-EQA. Note that this distribution resembles a Gaussian, with the median character count of queries being around , and word count being . This indicates that most generated queries in S-EQA are of an ideal length, with GPT-4 producing queries with concise scenario-specific language as opposed to an accumulation of multiple object states like MT-EQA [4].
IV-C Qualitative Results and Inference
In both our MTurk experiments (Refer III-C2), we measure average reviewer agreement across 5 users on each question to obtain ground truth labels on the data. For instance, if a question has Yes’s, and No’s, we label the ground truth as ‘Yes’. On the other hand, if even one of the annotations is ‘Cannot Answer’, we label the human-annotated output by the same.
For the first experiment, we obtain a low ‘Cannot Answer’ percentage of among the nearly participants. As such, a high percentage () of generated queries being answerable strongly indicates that the generated consensus object states are good representation of the human consensus for each query. Conversely, a low correlation of in LLM-generated and worker-annotated answers suggests the LLM’s limited capability in adequately answering them. To analyze this further, we attempt to reason about the generated answers by prompting GPT-4 to provide reasoning:
Situational Query and Object Data
Your answer is: LLM Answer (Yes/No).
Can you provide a brief reason for your answer focusing only on the object states and relationships provided?
Upon inspecting the generated reasoning, we notice that GPT-4 often goes against the consensus while predicting the answer. For example, consider the following datapoint -
Query: “Is the living room prepared for a movie night?”
Object-States : [tv: [‘ON’], lights: [‘OFF’], remotecontrol: [‘ON’], …]
Object-Relationships : [tv INSIDE livingroom, lights INSIDE livingroom, remotecontrol INSIDE livingroom, …]
LLM Answer: ‘NO’
The reasoning generated for the LLM Answer is as follows -
The lights are off which is suitable for a movie night, but the TV is on which is not suitable for movie watching as it may create distractions.
GPT-4 suggests that the TV being ON may create distractions for movie watching and hence responds with No. However, a more plausible situation would be that the user would watch the movie on the TV itself, and that the living room is indeed prepared for a movie night.
On the flip side, since our object states are limited to the ones present in VirtualHome, we are inclined to believe that the quality of the output depends on the richness of the object set. In this example, for instance, additional object information such as DVD: [‘Present’] and Popcorn ON Table would improve the consensus of a movie night about to happen. Since such objects do not exist in VirtualHome, it is reasonable to assume that GPT-4’s explanation is sufficient.
In our second MTurk survey, we sought to assess the usability of S-EQA by comparing situational and simple queries. For simple queries, of participants answer accurately, indicating their mostly unambiguous and straightforward nature. In contrast, S-EQA queries show only a accuracy. of users also choose Cannot Answer on the situational queries, as opposed to on the simple ones. This discrepancy showcases the usability of S-EQA in providing consensus objects and states for indirectly answering situational queries.
We infer from these qualitative experiments that while GPT-4 is indeed capable of producing situational queries and authentic consensus information, it often goes against the norm in answering the query itself. As such, an agent equipped with an LLM for directly answering situational queries from scene images might not be able to do so correctly. S-EQA is very useful here, as it simplifies the QA task by providing authentic objects and states relevant to the situational query for the agent to ground and subsequently answer the query.
V Conclusion, Limitations and Future Work
We tackle the novel and challenging problem of Embodied Question Answering with Situational Queries. Unlike prior work that dealt with simple objective queries pertaining target objects, situational queries necessitate an approximation of a consensus on object states. We use a Large Language Model (LLM), GPT-4 to approximate this consensus, incorporating it in an approach to pose unique situational queries, corresponding consensus object information and predicted answers. We use VirtualHome as an embodied simulation platform for our experiments, as it provides us with modifiable object states that also visually change in their appearance upon modification, allowing for quantitative validation.
We validate the authenticity of the generated data via a large scale MTurk user study with participants, and introduce it as S-EQA, the first EQA dataset addressing situational queries. In the study, we observe a low correlation of the GPT-4’s answers with that of the user, and notice a very high percentage of users () believe the queries are answerable with the provided consensus states. This lets us infer that while GPT-4 is capable of reliably generating consensus objects and states required for answering a situational query, it is incapable of directly answering the situational query itself. From this, we infer S-EQA’s usability, in providing a human-verified dataset of consensus data that could be used as a prior to break down situational queries into simpler, objective ones based on a human-consensus to answer them in household environments. Further, we perform Room and Object level VQA on S-EQA and quantitatively observe this discrepancy, and setting a benchmark for validating S-EQA on VirtualHome.
As a future experiment, we will look at a real-world deployment of an S-EQA-powered household robot, to gather data and finetune the LLM’s generational capabilities. The queries in S-EQA are limited due to the abstract, situational nature of the questions being generated and the limited number of objects. Future work will explore cases when the user poses novel questions outside the S-EQA dataset, requiring semantic alignment of the novel query with those in the dataset to extract consensus objects and states for S-EQA.

### v3

@@
-S-EQA: Tackling Situational Queries in Embodied Question Answering
+Is the House Ready For Sleeptime? Generating and Evaluating Situational Queries for Embodied Question Answering
 Abstract
 We present and tackle the problem of Embodied Question Answering (EQA) with Situational Queries (S-EQA) in a household environment.
-Unlike prior EQA work tackling simple queries that directly reference target objects and quantifiable properties pertaining them, EQA with situational queries (such as “Is the bathroom clean and dry?”) is more challenging, as the agent needs to figure out not just what the target objects pertaining to the query are, but also requires a consensus on their states to be answerable.
-Towards this objective, we first introduce a novel Prompt-Generate-Evaluate (PGE) scheme that wraps around an LLM’s output to create a dataset of unique situational queries, corresponding consensus object information, and predicted answers.
-PGE maintains uniqueness among the generated queries, using multiple forms of semantic similarity.
-We validate the generated dataset via a large scale user-study conducted on M-Turk, and introduce it as S-EQA, the first dataset tackling EQA with situational queries.
-Our user study establishes the authenticity of S-EQA with a high of the generated queries being deemed answerable, given the consensus object data.
-Conversely, we observe a low correlation of on the LLM-predicted answers to human-evaluated ones; indicating the LLM’s poor capability in directly answering situational queries, while establishing S-EQA’s usability in providing a human-validated consensus for an indirect solution.
-We evaluate S-EQA via Visual Question Answering (VQA) on VirtualHome, which unlike other simulators, contains several objects with modifiable states that also visually appear different upon modification — enabling us to set a quantitative benchmark for S-EQA.
-To the best of our knowledge, this is the first work to introduce EQA with situational queries, and also the first to use a generative approach for query creation.
-Our work introduces a new task and approach that aims to improve the usability of embodied agents in real-world household environments.
+Unlike prior EQA work tackling simple queries that directly reference target objects and properties (“What is the color of the car?”), situational queries (such as “Is the house ready for sleeptime?”) are challenging as they require the agent to correctly identify multiple object-states (Doors: Closed, Lights: Off, etc.) and reach a consensus on their states for an answer.
+Towards this objective, we first introduce a novel Prompt-Generate-Evaluate (PGE) scheme that wraps around an LLM’s output to generate unique situational queries and corresponding consensus object information.
+PGE is used to generate datapoints in the VirtualHome simulator, which is then annotated for ground truth answers via a large scale user-study conducted on M-Turk.
+With a high rate of answerability () on this study, we establish that LLMs are good at generating situational data.
+However, in evaluating the data using an LLM, we observe a low correlation of with the ground truth human annotations; indicating that while LLMs are good at generating situational data, they struggle to answer them according to consensus.
+When asked for reasoning, we observe the LLM often goes against commonsense in justifying its answer.
+Finally, we utilize PGE to generate situational data in a real-world environment, exposing LLM hallucination in generating reliable object-states when a structured scene graph is unavailable.
+To the best of our knowledge, this is the first work to introduce EQA in the context of situational queries and also the first to present a generative approach for query creation.
+We aim to foster research on improving the real-world usability of embodied agents through this work.
 I Introduction
@@
 A household robot deployed in the real world is likely to encounter such queries that require it to utilize consensus knowledge pertaining which objects and associated states are commonly associated with the situation.
-We term these as situational queries, and in this work, tackle the twofold challenge of 1) conjuring authentic situational queries and corresponding consensus object data and 2) highlighting the role of the consensus in simplifying the process of responding to situational queries.
+We term these as situational queries, and in this work, tackle the twofold challenge of 1) conjuring authentic situational queries and corresponding consensus object data and 2) highlighting the role of consensus object states in simplifying the process of responding to situational queries.
 A query such as “Is the house ready for sleeptime?” is a situational query, since there exist generally accepted object states, such as dimmed lights, closed windows, locked doors, etc., which most people associate with a house being ready for nighttime.
-However, a query like “Can I bring guests over?”, while situational, is too ambiguous, as the answer relies on various unstated factors and user preferences on household norms, and event-specific constraints.
-For situational queries to be validatable, it is important for them to not be too ambiguous so as to allow for a consensus to exist.
-Our work places emphasis on this, and follows an in-context learning scheme on an LLM to avoid overly simplistic nor excessively ambiguous queries.
-Developing a dataset for EQA in the past has been limited to constructing simple queries using entities from a simulator [4, 5], with some work extending it to abstract queries built on scene-graphs [6].
+However, a query like “Can I bring guests over?”, while situational, is too personal, as the answer relies on various user preferences on household norms and event-specific constraints.
+In our work, we aim to generate and human-validate queries where a general consensus exists on multiple object states.
+Developing EQA datasets in the past has been limited to constructing simple queries using entities from a simulator [5, 6], with some work extending it to abstract queries built on scene-graphs [7].
 Posing situational queries using these existing techniques is however nontrivial, as they stem from human intuition and contextual awareness of the environment.
-Setting up real-world experiments to collect this data is also cumbersome; if not having an intrusive agent capturing conversations at home, an offline approach would be cognitively heavy, requiring humans to come up with queries for imaginary situations.
-As such, in our work, we use an LLM to generate and study such queries, while ensuring they cover unique concepts stemming from various situations.
-Table I compares our dataset with existing ones.
-Main Results: To tackle these issues, we leverage the commonsense capabilities of Large Language Models (LLMs) and present an approach to generate situational queries and associated consensus object states and relationships.
-LLMs allow us to generate situational query data for a variety of scenarios which are nontrivial to obtain via real-world data collection.
+Setting up real-world experiments to collect this data is also cumbersome; if not having an intrusive agent capturing conversations at home, an offline approach would be cognitively heavy, requiring humans to imagine situations for queries.
+In our work, we develop a scheme that utilizes a Large Language Model (LLM) to generate such queries, providing in-context cures to cover unique concepts stemming from various situations.
+Table I compares our dataset with existing work.
+Main Results: We tackle the issues presented above, utilizing LLMs as a prior for generating situational queries and associated consensus object states and relationships111https://youtu.be/m3hg59zGYg8.
+We human-validate the generated data, and present a benchmark for EQA on situational queries.
 The main contributions of our work are as follows:-
-We propose a novel Large Language Model (LLM)-based method for generating Situational Queries for Embodied Question Answering (EQA).
-Unlike simple queries that directly reference particular objects, situational queries necessitate an understanding of the consensus objects and states pertaining a situation.
-Our proposed Prompt-Generate-Evaluate (PGE) scheme wraps around an LLM to generate a wide array of unique and challenging queries covering various household scenarios.
-Along with the queries, we also generate corresponding consensus object information, pointing to object states and relationships relevant to the queries.
-The generated queries are validated via a large scale MTurk [7] user study to produce S-EQA, a LLM-generated dataset containing situational queries, associated consensus object data, and predicted answers.
-While the number of queries is small due to their highly abstract nature, PGE ensures diversity in the queries across rooms.
-We quantitatively evaluate S-EQA by performing Visual Question Answering (VQA) on VirtualHome images with both situational queries (Room VQA) and consensus data (Object VQA) to establish a benchmark.
-We observe that 97.26% of users find the generated queries to be answerable provided the consensus data, while only 46.2% of user answers align with the LLM predicted ones.
-From the former, we can infer the authenticity of LLM generated data in producing human-like consensus states, while the latter indicates the LLM cannot be used to directly answer situational queries; establishing the need and use-case for S-EQA, a human-validated dataset that provides situational queries along with consensus objects and states.
+Problem: We tackle the novel problem of posing and evaluating Situational Queries for Embodied Question Answering (S-EQA).
+Unlike simple queries that directly reference particular objects, posing situational queries itself is challenging, requiring a consensus on multiple objects and their states, and cannot be formed easily using prior rule-based methods (Refer Table I).
+Method: We propose a novel Prompt-Generate-Evaluate (PGE) scheme that integrates an LLM with a feedback scheme that allows for the generation of several unique situational queries, along with corresponding consensus object-state and object-relationship information that needs to be satisfied to successfully answer the question.
+Human-Validation: Using PGE in the VirtualHome simulator, we produce a S-EQA dataset of situational queries and consensus data.
+We annotate these via a large scale MTurk [8] user study, which gives us a high authenticity percentage of , indicating that the data generated corresponds to consensus.
+LLM Eval: We prompt the LLM to provide binary answers to S-EQA datapoints generated from VirtualHome.
+Despite being able to generate highly authentic situational data, we observe a subpar correlation of with the human annotations when it comes to answering.
+We infer that LLMs can generate good situational queries and consensus object-states, but answer them poorly.
+LLM Reasoning: When asked for an explanation, we observe the poor commonsense reasoning of LLMs in justifying their answer. This further suggests that LLMs exhibit poor situational awareness.
+Real World Challenges: Finally, we discuss the transfer of PGE to a real world real world lab environment, and discuss various challenges with LLM hallucination that obstruct the generation of quality consensus data.
 II Related Work
@@
 However, these queries are still limited to simple object grounding within a scene, and do not require any form of situational awareness about the readiness of an environment.
-For instance, S-EQA considers queries like “Is the living room ready for a movie night?”, and presents consensus multiple object states (TV: On, Popcorn: Cooked, On Table, etc.) to measure this readiness.
-For the same example, a K-EQA query would have been constructed as “Is there a device for watching movies in the living room?”, which requires only scene knowledge for an answer, but does not require situational knowledge about the consensus states of objects for a movie night.
-In our work, we utilize an LLM to generate situational queries and approximate human consensus on object states required to answer the query.
+For instance, S-EQA considers queries like “Is the living room ready for a movie night?”, that requires consensus on multiple objects and their states (TV: On, Popcorn: Cooked, On Table, etc.) to measure this readiness.
+A K-EQA query for the same example could be constructed as “Is there a device for watching movies in the living room?”, which requires only scene knowledge about the availability of a TV, but does not require situational knowledge about the various object-states involved in a movie night.
+In our work, we utilize an LLM to generate situational queries and approximate human consensus on object states.
 Our queries are an abstraction above the scene-knowledge based K-EQA queries, requiring consensus knowledge about multiple object states for validation.
@@
 In embodied AI, these capabilities have recently been leveraged to produce superior planning and decision-making strategies during navigation [18, 19, 20, 21].
-In our case, we wish to generate situational queries for embodied question answering, which as defined in Section I require consensus about object states and relationships in the environment.
-As such, we are motivated to use LLMs for generating these queries and ask if their commonsense reasoning capabilities can sufficiently represent the consensus needed to answer situational queries.
-III Situational EQA: Generation, Analysis, and Evaluation
+In our case, we wish to generate situational queries for embodied question answering and ask if the common sense reasoning capabilities of LLMs can both generate such queries as well as the the consensus object-states needed for sufficiently answering them.
+III Situational EQA: Generation & Validation
 Having motivated the problem of Situational EQA, we now present a detailed description of our approach.
-We present three phases — generating situational data using LLMs, validating this data via segregation and a user survey, and finally evaluating it on the VirtualHome simulator.
-We hope the generative approach being used to develop S-EQA helps foster future research in this domain.
+We present three phases — generating situational data using an LLM (GPT-4), annotating the data via a large-scale user survey to produce a S-EQA dataset, and finally evaluating it on the VirtualHome simulator.
 III-A Situational Query Definition
@@
 are in their appropriate state.
-Additionally, we want to ensure that the queries are related to a situation and are not simple in nature, i.e., the target objects and states are not in the query itself.
+Additionally, we want to ensure that the queries are related to a situation and are not simple in nature, i.e., the target objects and states are not referenced in the query itself.
 For example, “Is the sofa blue?” [2] or “How many objects can be used to hold rice?” [7] are simple queries, as they do not deal with a situation, but rather pose a query referencing objects in the environment.
@@
 Considering the LLM as a function that generates queries , and to be a set of all possible queries generated by , we want situational queries , which satisfies the following conditions -
-The query must not directly reference any object, i.e., it must omit words from a set which includes “object”, “table”, “chair”, etc.
-The query must be a question.
-The answer to the query must be binary i.e., Yes/No, and not subjective in nature.
+Abstraction: The query must not directly reference any object, i.e., it must omit words from a set containing words such as “object”, “table”, “chair”, etc.
+Contextual: The query requires an assessment of relevant objects and their states collectively satisfy a scenario (e.g., “Is the dining area set up for dinner?”), requiring identification of appropriate objects and consensus on their expected states (e.g., table present, napkins, lighting, etc.).
+Binary: The answer to the query must be binary i.e., Yes/No, and not subjective in nature.
 Formally, we can define as:
@@
 III-B Generating Situational Queries: S-EQA
-Figure 2 presents an overview of our generation approach.
-We generate datapoints of the format - [Situational Query, Object-States, Object-Relationships, LLM Answer].
-A sample datapoint is as follows -
-Query: “Is the bathroom ready for a shower?”
+We prompt the LLM to generate datapoints of the format - [Situational Query, Object-States, Object-Relationships]. For example,
+Situational Query:
+“Is the bathroom ready for a shower?”
 Object-States : [lightswitch: [‘On’], towels: [‘Present’], Soap: [‘Present’] …]
 Object-Relationships : [lightswitch inside bathroom, towels inside bathroom, … ]
-LLM Answer: ‘Yes’
-Here, the LLM points to a set of object states and relationships (such as ‘towels inside bathroom being present’) that it believes need to be checked to determine if the bathroom is ready for a shower.
-We also predict an answer to the query that is later verified in our user study.
-III-B1 VirtualHome Setup
-We use the VirtualHome [21] simulator for defining ground truth objects, states and properties in our experiments.
+The object states and relationships (such as ‘towels inside bathroom being present’) being generated represent scene-graph requirements that the LLM believes need to be checked for the query to satisfy equation 1.
+VirtualHome Setup: We use the VirtualHome [1] simulator for defining ground truth objects, states and properties in our experiments.
 It contains 7 different household environments containing over 390 objects.
@@
 We use these room markings to group and analyze the generated queries later on.
-III-B2 Prompt-Generate-Evaluate (PGE) Scheme
-Figure 2 presents an overview of our Prompt-Generate-Evaluate scheme for generating data.
-We perform analytical operations on the LLM’s generated datapoints to ensure they fit into our definition of situational queries, while also maintaining diversity.
-The System Prompt for our LLM contains a list of all object states and relationships.
+Prompt-Generate-Evaluate (PGE) Scheme: Figure 2 presents an overview of our Prompt-Generate-Evaluate scheme for generating data.
+We analyze the LLM’s generated datapoints to provide feedback for ensuring the data fits into our definition of situational queries, while also maintaining diversity.
+The System Prompt for our LLM contains a list of all object states and relationships, which in this case is the scene-graph provided by VirtualHome.
 We also provide additional rules and in-context examples to improve the generation of prompts.
-For instance, explicitly mentioning “Do not conjure any new object states” ensures that the generated datapoints do not contain states that are not achievable in VirtualHome.
+For instance, explicitly mentioning “Do not conjure any new objects or states” ensures that the generated data does not contain objects not present or states that are not achievable in VirtualHome.
 Additionally, we also provide a set of “Generated Queries” as feedback to the system prompt.
-These represent queries that the LLM has already generated, and in the User prompt, encourage the LLM to generate queries different from these.
-Our strategy for choosing these queries is discussed in the feedback measures presented below.
-Our User Prompt asks to generate datapoints in the format specified earlier, and provides a few sample datapoints.
-We also provide some in-context rules to ensure that the generated queries are not too simplistic, not too ambiguous, and are also unique.
+These queries are representative of what the LLM has already generated, and we encourage the LLM to generate queries different from these.
+Our User Prompt asks to generate datapoints in the format specified earlier, and while providing a few manually annotated sample datapoints.
+Additionally, we provide some in-context rules to ensure that the generated queries are not too simplistic, not too ambiguous, and are also unique.
 For instance, we explicitly state ‘The situational query must not reference any object or even contain the word ‘object’.’ and ‘The query must have a Yes/No answer.’.
-Additionally, we ask it to generate queries very different from what is already present in the Generated Queries in the System Prompt, to avoid simple sentence rephrasing.
-To ensure that our Situational Query Database contains data points that adhere to the specifications presented above, we take the following two feedback measures:
-After generating data points, we embed the situational queries using RoBERTa [22] to get sentence embeddings .
+We also ask it to generate queries very different from what is already present in the Generated Queries in the System Prompt and come up with ‘unique and creative household situations‘ with the object data provided.
+To ensure that our Situational Query Database contains diverse data points adhering to specifications presented in equation 1, we take the following two feedback measures:
+Generated Query Update: After generating data points, we embed the situational queries using RoBERTa [22] to get sentence embeddings .
 We perform agglomerative clustering [23] on to accumulate them into groups of queries based on sentence similarity.
@@
 Finally, we update the system prompt with these queries.
-Simultaneously, we compute the cosine similarity of the generated query embeddings with the queries already present in the Situational Query Database .
-If is greater than a particular threshold (), we continue the “conversation” with the LLM, by presenting it with a re-generation prompt.
+Continued Conversation: Simultaneously, we compute the cosine similarity of the generated query embeddings with the queries already present in the Situational Query Database .
+If is greater than a particular threshold (), we continue the conversation with the LLM, by presenting it with a re-generation prompt.
 We mention the percentage of similarity, and ask the LLM to re-generate a new batch of datapoints.
 These feedback measures ensure contains unique situational queries that are not simple rewordings, while also encouraging diversity in the generated situations.
+Contrasting examples with and without PGE are illustrated in Fig.
+4.
 The total number of datapoints generated is always less than or equal to the product of the looping parameters and , i.e., .
@@
 We validate this data via a large-scale MTurk user study, and present this as S-EQA, a dataset containing diverse and challenging situational queries along with consensus object data to help answer them.
-III-C Validating S-EQA: t-SNE Plots and MTurk Survey
-Our objective with validating S-EQA is not just to measure the authenticity of the generated data, but also to establish necessity and the real-world usability of our dataset.
-First, we first visualize the queries in using t-SNE plots to analyze their diversity.
-We then conduct a Mechanical Turk based user study to validate the generated data.
-III-C1 Analyzing S-EQA with t-SNE
-Figure 3 illustrates t-SNE plots we derive from . These are as follows -
-Room Categorization: Using the room properties obtained from VirtualHome (section III-B1), we group the questions into 6 categories - [kitchen, livingroom, bedroom, bathroom, multi-room, no room] (Ref.
-Fig.
-3(a)).
+Prompts: We use the following prompts for PGE.
+System Prompt:
+“I have a list of objects, and their states, and relationships in a household.
+Object states are listed in the OBJ_STATE_DICT dictionary below. Each item has the format OBJECT: [STATES].
+STATES, IF PRESENT, can switch between ON/OFF and OPEN/CLOSED. Do not conjure any nrew objects or states.
+OBJ_REL_DICT contains the initial relationships between the objects. These can be changed. For instance, "apple INSIDE fridge" is a valid relation.
+OBJ_STATE_DICT :- {…}
+OBJ_REL_DICT :- {…}
+GENERATED_QUERIES :- {…} ”
+User Prompt:
+“ Using OBJ_STATE_DICT and OBJ_REL_DICT, can you generate {X} potential questions, states and relationships that the user might ask about the environment?
+These must be about a potential scenario, requiring situational awareness and a consensus on multiple object, their states and relationships.
+The output should be lines of the form [Question, Object-State Pairs, Relationships].
+Make sure you generate questions very different from those in GENERATED_QUERIES. Get creative with the potential scenarios!
+Make sure to use an exhaustive set of relationships and object-state pairs for each query.
+The query must have a Yes/No answer.
+The query must not directly reference any object or even contain the word ‘object’. ”
+Re-Generation Prompt:
+“{X}% of the questions are similar to what you’ve already generated earlier! Try again, give me {Y} more.
+QUERIES:…”
+III-C Validating S-EQA: t-SNE plots and M-Turk Annotation
+Our objective with validating S-EQA is not just to annotate the queries, but to also measure the authenticity of the generated data.
+We first visualize the queries in using t-SNE plots to analyze their diversity.
+We then conduct a user study via Mechanical Turk [8] to annotate the generated data.
+t-SNE Plots: We embed the generated queries with BERT and visualize them using t-SNE.
+They are then are categorized by Room (pick from [kitchen, living room, bedroom, bathroom, multi-room, no room]), Situational/Simple (Yes/No) and Spatio-Temporal (time-sensitivity).
+We observe that GPT-4 tends to generate queries specific to objects in a single room over multi-room or no-room ones like “Is the house ready for sleeptime?”.
+Further, we note that most queries () are classified as situational and not simple.
+Around are categorized as spatial queries (like “Is the kitchen ready for cooking?”) that refer to the current environment state, as opposed queries like “Was someone in the kitchen today?” which require deeper situational understanding involving time.
+Overall, we observe that most queries adhere to the definition provided in Equation 1.
+Room Categorization: Using the room properties obtained from VirtualHome (section III-B), we group the questions into 6 categories - [kitchen, livingroom, bedroom, bathroom, multi-room, no room].
 We infer from this plot that most questions being generated belong to one particular room, as opposed to multiple ones.
@@
 We prompt and present in-context examples accordingly.
-We notice that of generated questions (Ref.
-Fig.
-3(b)) are classified as situational according to the LLM.
+We notice that of generated questions are classified as situational according to the LLM.
 As the LLM classifications have several false positives, this result motivates us to use human-validation to authenticate the dataset.
@@
 We observe that of the generated questions are Spatial ones, showing that PGE prefers modeling spatial relations over temporal ones.
-III-C2 Assessing S-EQA’s Utility with MTurk
-We conduct a large-scale Mechanical Turk [7] study to gauge S-EQA’s authenticity and usability in answering real-world queries.
-In our first experiment, we assess whether the consensus object data being generated adequately facilitates answering situational queries.
-Workers are shown a query, images of the corresponding room, and images of objects set to consensus states, while the LLM Answer remains hidden to negate bias.
+MTurk Validation: We conduct a large-scale Mechanical Turk [8] study to annotate the generated data and evaluate the authenticity of real-world queries.
+The annotators are chosen from the US-East region, with at least a approval rate on previous work.
+Annotators are shown a generated query, along with consensus object-states and relationships.
 Given options of Yes, No, and Cannot Answer, a worker’s capacity to give a definitive response, i.e., choosing Yes or No, serves as an indicator of whether the generated data is in alignment with general user consensus.
-For instance, the query “Is the house ready for sleeptime?”, if unanimously agreed upon by workers when shown images of dimly-lit rooms, closed doors, and windows, validates the data point.
-Subsequently, in a second experiment, workers are presented with an image of a room accompanied by a situational S-EQA query (For eg.
+For instance, the query “Is the house ready for sleeptime?”, if agreed upon by workers when shown images of dimly-lit rooms, closed doors, and windows, validates the data point.
+We obtain a low ‘Cannot Answer’ percentage of among the participants.
+As such, a high percentage of generated queries being answerable strongly indicates that the LLM is capable of generating object-states and queries that are a good representation of human consensus.
+Subsequently, we conduct a second comparison study with the same workers, where they are presented with an image of a room accompanied by a situational S-EQA query (For eg.
 Is someone working in the bedroom?
-with top-views of the bedroom) and, separately, images of consensus objects paired with simpler, objective queries about their state (For eg., Is the Computer On?
+with top-views of the bedroom) and, separately, images of objects paired with consensus queries formulated based on their state (For eg., Is the Computer On?
 with an image of a computer being on).
-For the simple query annotation, we additionally sample ‘Yes/No’ questions the SimVQA [24] dataset to even out the number of situational and simple queries.
-An improved annotator performance in simple over situational queries would indicate S-EQA’s real-world usability in providing human-authenticated consensus object-states for situational queries which are relatively simpler to answer.
-III-D Evaluating S-EQA in VirtualHome
-Assuming low-level navigation to be solved, we quantitatively evaluate S-EQA via Visual Question Answering (VQA).
-VirtualHome provides us with a household environment where object states visually change upon modifying states.
-We present VQA models with the same setup as described in our first MTurk experiment presented in III-C2, and compare the results.
-Figure 4 presents an overview.
+We sample additional ‘Yes/No’ questions the SimVQA [24] dataset to even out the number of situational and consensus queries.
+An improved annotator performance in answering consensus queries over situational ones would indicate S-EQA’s utility in breaking down subjective queries into simpler consensus queries for improved question-answering.
+We discuss this later in the results section.
+In both our MTurk experiments, we measure average reviewer agreement across 5 users on each question to obtain ground truth labels on the data.
+For instance, if a question has Yes’s, and No’s, we label the ground truth as ‘Yes’.
+On the other hand, if even one of the annotations is ‘Cannot Answer’, we label the human-annotated output by the same.
+IV Situational EQA: Evaluation
+Figure 5 presents an overview of our evaluation of the simulator-based dataset, and Figure 6 presents our real-world setup for evaluation.
+We evaluate queries both via VQA as well as by using an LLM.
+Further, LLM answers are qualitatively studied via reasoning to infer situational awareness.
+IV-A VQA Evaluation:
+We perform VQA on images taken from VirtualHome with situational and consensus queries.
+The VirtualHome simulator is unique in enabling this, as it allows us to modify consensus object states pertaining to subjective queries such that they also visually change upon modifying states.
+Situational Consensus Queries: Consensus object states associated with the generated situational queries are used to formulate binary objective questions for evaluation. These are of the form -
+Is the OBJECT STATE?
+For instance, for a query ‘Was someone working in the bedroom?’, the consensus queries would be ‘Is the Computer On?’ and ‘Is the Lightswitch On?’.
+We present VQA models with both situational and consensus queries, and compare the results.
 Room VQA (Situational Query): Here, we present the VQA model with multiple top-view images of a relevant room as input, and ask it to answer the associated situational query with Yes/No/Cannot Answer.
-Most of our queries are associated with one particular room (see Figure 3(a)), which is often part of the query itself.
-During our experimentation, we observed a clear bias in model performance with such queries.
+The assumption here is that the agent might have access to overhead security cameras which it could use to answer the queries.
+Most of our queries are associated with one particular room, which is often part of the query itself.
+During our experimentation, we observe a bias in model performance when the name of the room itself is present in the query.
 For instance, if the query contained ‘living room’, and the image was that of the living room, the answer would always be yes, irrespective of the actual context of the query.
 To mitigate this, we replace any mention of a room in the query with the generic term - “this place”.
-Object VQA (Simple Query): Here, we use consensus object states linked with the situational query to formulate simple objective questions for a VQA model. These are of the form -
-Is the OBJECT STATE?
-We provide the VQA model with questions along with images of the consensus objects in their designated states.
+Object VQA (Consensus Query): We provide the VQA model with these consensus queries along with images of the corresponding objects in their designated states.
 For both Room VQA and Object VQA, we measure the accuracy and F1-scores of prediction across all the images, using the human-verified answers as ground truth.
@@
 is then defined as:
-and give us an upper bound on the performance of VQA models on S-EQA in VirtualHome. For , we set a threshold of above which the score would be deemed satisfactory.
-IV Results and Inferences
-We use GPT-4 [12] as the LLM of choice to generate situational queries in our experiments.
-Our ablations with GPT-3.5 showed poor quality in terms of generation, with many outputs not fitting into our specified datapoint format.
-IV-A Preparing the Simulator
+and give us an upper bound on the performance of VQA models. For , we set a threshold of above which the score would be deemed satisfactory.
+IV-B LLM Evaluation:
+Motivated by K-EQA [7], we utilize scene graphs provided by VirtualHome to develop an LLM-based action planner.
+For S-EQA, we have situational queries generated along with consensus object states and relationships.
+VirtualHome provides us with a fully observable scene graph for this, i.e., the states and relationships of all objects are known.
+Our aim is to measure how well the LLM is able to approximate the human consensus behind answering situational queries.
+For this, we first ask the LLM to provide a Yes/No answer to each of the the datapoints generated.
+We then measure the correlation between the LLM and human annotated answer.
+Further, we also ask the LLM to reason about its answers for a subjective analysis.
+IV-C Real World Experiments:
+Unlike VirtualHome, where a fully observable scene graph is readily available, the real-world environment does not have object states in a predefined format.
+Instead, we gather images from various locations in our environment which is used as context for GPT-4V to generate situational datapoints.
+A representative subset of scenes and datapoints being generated are illustrated in Fig 6.
+We aimed to create diverse situations including group meetings, robot operations, cafeteria related events across various locations in our lab environment.
+The VLM predicts Yes/No responses for each datapoint.
+These are then compared against human-annotated ground truth responses to assess alignment.
+We then analyze failure cases where the LLM misinterprets missing information or hallucinates object states, showcasing the difficulty in accurately comprehending situational queries in the real world.
+V Results and Inferences
+GPT-4 [13] is our LLM of choice to both generate and evaluate situational queries in our experiments.
+Our ablations with GPT-3.5 showed poor quality in terms of generation, with many outputs not fitting the specified data format.
+V-A Simulator Setup
 We first program the object states and relationships from the generated data points into the VirtualHome environment.
-This also includes filtering out a small subset of bad states which are not feasible.
+This also includes manually filtering out a small subset of bad states which are not feasible.
 For instance, ‘computer INSIDE fridge’ is not a valid relationship for us to set, and the simulator prevents this.
-For VQA and the MTurk user study, we require images of the room and corresponding consensus objects set to their appropriate states.
+For VQA evaluation, we require images of the room and corresponding consensus objects set to their appropriate states.
 We gather room images by considering the 4 corners of each room, with a camera pointing downwards towards the center.
@@
 Then for object images, we use the first-person view of a human placed in the environment.
-The human is instructed to walk to each of the objects present in a S-EQA datapoint, and we grab an image when they are present in front of them.
-This approach for capturing images proved simpler to implement and is in-line with the visual perspective of an agent grabbing an object.
-IV-B Quantitative Results and Benchmark
-We use BLIP-2 [25], BLIP [26] and ViLT [27], all finetuned on VQAv2 [28] to evaluate S-EQA, and these results are presented in Table II.
+We consider the human to be our robot agent here.
+They are instructed to walk to each of the objects present in a S-EQA datapoint, where we grab an image when they are present in front of them.
+This approach for capturing images proved to be simpler to implement over simulator camera placement and gives us a visual perspective of an agent grabbing an object.
+V-B VQA Results
 | VQA Model | Accuracy % | F1-Score % | |
@@
 | ViLT [27] |
-Observe the higher Object-VQA accuracies and F1 scores over Room-VQA across each approach.
-This suggests that answering situational queries by looking at images at a room level is significantly harder than answering simple queries about the states of target objects.
-Our second MTurk experiment (Refer III-C2) further corroborates this result, suggesting S-EQA’s usability in providing the simpler consensus object data for answering situational queries.
+Observe the higher Object-VQA accuracies ( on average) and F1 scores over Room-VQA across each approach.
+This suggests that answering situational queries by looking at images at a room level is significantly harder than answering queries about the states of target objects.
+Our comparison study with MTurk (Refer III-C) further corroborates this result, with of users finding consensus queries to be simpler to answer over situational ones.
+This suggests that our approach of breaking down situational queries into simpler consensus ones could be used to indirectly answer difficult situational queries.
 The high values on joint scores indicate an upper bound on performance.
-While still better than RoomVQA, ObjectVQA models have mediocre performance overall, despite the queries being simple.
-This can be attributed to inference on synthetic images from VirtualHome, which these models have not been trained on.
-Our real-world experiments suggest improved performance, confirming this domain gap.
+While still better than RoomVQA, ObjectVQA models have mediocre performance overall, despite the queries being straightforward in nature.
+This can be attributed to inference on synthetic images from VirtualHome which these models have not been trained on, opening an avenue for future work.
+Language - Question length, re-classification percentage, Plot curve of average length of dataset questions. VQA - Room VQA and Object VQA
 Figure 7 provides a distribution of the query lengths in S-EQA.
@@
 This indicates that most generated queries in S-EQA are of an ideal length, with GPT-4 producing queries with concise scenario-specific language as opposed to an accumulation of multiple object states like MT-EQA [5].
-IV-C Qualitative Results and Inference
-In both our MTurk experiments (Refer III-C2), we measure average reviewer agreement across 5 users on each question to obtain ground truth labels on the data.
-For instance, if a question has Yes’s, and No’s, we label the ground truth as ‘Yes’.
-On the other hand, if even one of the annotations is ‘Cannot Answer’, we label the human-annotated output by the same.
-For the first experiment, we obtain a low ‘Cannot Answer’ percentage of among the nearly participants.
-As such, a high percentage () of generated queries being answerable strongly indicates that the generated consensus object states are good representation of the human consensus for each query.
-Conversely, a low correlation of in LLM-generated and worker-annotated answers suggests the LLM’s limited capability in adequately answering them.
-To analyze this further, we attempt to reason about the generated answers by prompting GPT-4 to provide reasoning:
+V-C LLM Eval Results
+We obtain a low correlation of between GPT-4 generated answers and the human-annotated ones.
+This is despite the LLM having complete access to the entire scene graph which it used earlier to generate situational queries.
+Furthermore, the low ‘Cannot Answer’ percentage of on the annotated responses shows that most queries and consensus data generated did indeed make sense to most users.
+We can infer from these two values that while GPT-4 is good at generating situational queries and the potential consensus data needed to answer them, it is poor at answering it’s own generated queries.
+To understand this further, we prompt for reasoning as follows:-
+LLM Reasoning: To analyze the poor LLM performance on query answering, we reason about the generated answers by prompting GPT-4 to provide reasoning as follows:
 Situational Query and Object Data
@@
 LLM Answer: ‘NO’
-The reasoning generated for the LLM Answer is as follows -
+The reasoning generated for the LLM Answer is -
 The lights are off which is suitable for a movie night, but the TV is on which is not suitable for movie watching as it may create distractions.
@@
 Since such objects do not exist in VirtualHome, it is reasonable to assume that GPT-4’s explanation is sufficient.
-In our second MTurk survey, we sought to assess the usability of S-EQA by comparing situational and simple queries.
-For simple queries, of participants answer accurately, indicating their mostly unambiguous and straightforward nature.
-In contrast, S-EQA queries show only a accuracy.
-of users also choose Cannot Answer on the situational queries, as opposed to on the simple ones.
-This discrepancy showcases the usability of S-EQA in providing consensus objects and states for indirectly answering situational queries.
-We infer from these qualitative experiments that while GPT-4 is indeed capable of producing situational queries and authentic consensus information, it often goes against the norm in answering the query itself.
-As such, an agent equipped with an LLM for directly answering situational queries from scene images might not be able to do so correctly.
-S-EQA is very useful here, as it simplifies the QA task by providing authentic objects and states relevant to the situational query for the agent to ground and subsequently answer the query.
-V Conclusion, Limitations and Future Work
-We tackle the novel and challenging problem of Embodied Question Answering with Situational Queries.
-Unlike prior work that dealt with simple objective queries pertaining target objects, situational queries necessitate an approximation of a consensus on object states.
-We use a Large Language Model (LLM), GPT-4 to approximate this consensus, incorporating it in an approach to pose unique situational queries, corresponding consensus object information and predicted answers.
-We use VirtualHome as an embodied simulation platform for our experiments, as it provides us with modifiable object states that also visually change in their appearance upon modification, allowing for quantitative validation.
-We validate the authenticity of the generated data via a large scale MTurk user study with participants, and introduce it as S-EQA, the first EQA dataset addressing situational queries.
-In the study, we observe a low correlation of the GPT-4’s answers with that of the user, and notice a very high percentage of users () believe the queries are answerable with the provided consensus states.
-This lets us infer that while GPT-4 is capable of reliably generating consensus objects and states required for answering a situational query, it is incapable of directly answering the situational query itself.
-From this, we infer S-EQA’s usability, in providing a human-verified dataset of consensus data that could be used as a prior to break down situational queries into simpler, objective ones based on a human-consensus to answer them in household environments.
-Further, we perform Room and Object level VQA on S-EQA and quantitatively observe this discrepancy, and setting a benchmark for validating S-EQA on VirtualHome.
-As a future experiment, we will look at a real-world deployment of an S-EQA-powered household robot, to gather data and finetune the LLM’s generational capabilities.
-The queries in S-EQA are limited due to the abstract, situational nature of the questions being generated and the limited number of objects.
-Future work will explore cases when the user poses novel questions outside the S-EQA dataset, requiring semantic alignment of the novel query with those in the dataset to extract consensus objects and states for S-EQA.
+These results suggest that while GPT-4 is indeed capable of producing situational queries and authentic consensus information, its displays poor commonsense reasoning while answering such queries.
+V-D Real World Experiments: Results and Discussion
+We experiment with situational queries in our real-world lab environment, assuming the presence of an assistive workplace robot agent.
+We gather a total of images across various situations, corresponding to different locations in our workplace, and feed these into PGE as visual context for GPT-4V.
+We then generate datapoints using PGE and manually annotate them.
+During annotation, we note several hallucinations in the object states that did not occur earlier when the scene graph was explicitly defined in VirtualHome.
+Some instances of this are highlighted in red in Figure 6.
+For example, note the noise level indication is present as an object state in many cases, but can not be discerned by looking at the image.
+Ghost objects such as microphones and safety equipment also appear in the generated object-states.
+We infer here that a grounded scene graph, like what VirtualHome provides is important to avoid poor data generation.
+Evaluating the generated datapoints with GPT-4 presents a mismatch of , a similar trend as that with the VirtualHome data, and solidifies our claim that LLM generated binary answers are misaligned with human consensus.
+Similar to what was done with VirtualHome, we conduct LLM based reasoning on the queries generated.
+An example of this is illustrated in Fig.
+8.
+Note the counterintuitive explanation that GPT gives when asked if the meeting space is ready for a group discussion.
+Despite the table being occupied with someone present in the room, who also appears to be using their laptop (Object State includes Laptop:OPEN), the reasoning is completely off.
+We observe many such cases in the data that we gather, where LLM answers and the justification provided goes against human consensus.
+LLMs have been widely adopted in robotics for their strong commonsense reasoning, enabling significant performance gains across embodied tasks [18, 16, 28].
+These models have contributed to the rise of foundation models for robotics, where a common paradigm involves LLMs predicting task outputs or structured action plans, which agents then execute with varying affordances [29].
+Given these successes, one would expect LLMs to exhibit similar strengths in our task.
+However, we observe a counterintuitive outcome—LLMs perform poorly in commonsense reasoning.
+While this may stem from the inherent subjectivity of situational queries, it also reveals GPT-4V’s limited situational awareness, even when provided with visual cues.
+Additionally, hallucinations in object states likely degrade downstream query performance, highlighting how even minor inaccuracies in state identification can significantly impact task execution.
+This contrasts with prior work where LLMs have demonstrated strong object reasoning, suggesting that situational awareness presents a distinct challenge.
+Future work will look at improving LLM-based situational reasoning by employing more robust multimodal alignment schemes via image captioning, memory-aware mechanisms, or explicit uncertainty modeling to mitigate state hallucinations and enhance task reliability.
+VI Conclusion, Limitations and Future Work
+We introduce and tackle the novel and challenging problem of Embodied Question Answering with Situational Queries — S-EQA.
+Unlike prior work that dealt with simple objective queries that directly reference target objects, situational queries necessitate an approximation of a consensus on multiple object states.
+Given the challenge in gathering such queries from the real world, we rely on commonsense knowledge from an LLM, GPT-4 to develop a Prompt-Generate-Evaluate (PGE) framework for posing unique situational queries and corresponding consensus object information.
+In the VirtualHome simulator, we use PGE to generate 2K unique queries, and annotate them via a large scale MTurk user study with participants.
+The high answerability () lets us infer that LLMs are good at generating situational queries and corresponding consensus data.
+However, using GPT-4 to answer the simulator queries shows a low correlation with the human annotations (); we infer that LLMs are good at generating situational data but poor at answering them.
+When asked to explain its answers, we observe the LLM’s poor commonsense capability, exhibiting counterintuitive reasoning.
+This suggests that LLMs show poor situational awareness.
+Finally, we discuss challenges with transferring our generation and evaluation schemes to a real world lab environment.
+Since a grounded scene graph isn’t directly available for use in the real-world, we provide images of various locations as context for query generation, and observe several unintended hallucinations with ghost object states as consensus.
+Future work will focus on limiting these by first gathering structured scene graphs using existing visual map building techniques.